OpenStudy (anonymous):

can some one help me? Parallel Lines "L" and "M" are cut by a transversal. Angle "A" measures (15 x + 6 ) degrees and angle "F" measures (5x+14) degrees. What are the degree measures of angles c & h ? Please try to show your work.

OpenStudy (anonymous):

would help if we had an image with angles labeled if you post one i will help

OpenStudy (anonymous):

ignore the bottom part. that has nothing to do with my math. top parallel line is line "L" and bottom parallel line is line "M"

OpenStudy (anonymous):

ive tried.... 15x + 6 = 5x + 14 -5x -5x 10x + 6 = 14 - 6 - 6 10x = 8 and it makes no sense. i need to find the angles of C and H

OpenStudy (anonymous):

A+C=180, A=H, and F=c. this stems from the fact that the two lines being intersected are parallel. susbstitution gives us A+F=180. plugging in the given values for A and F you get (15x+6)+(5x+14)=180 which reduced is 20x+20=180. from there it is just algebra to find x and find the numerical values of H and C